OpenStudy (anonymous):

Help please :) Paul is deciding whether to buy a new television set with special store financing. The financing order is 2% simple interest for 5 years. The television paul wants cost $650. What would be the total he would be paying the store if he takes the special finance after? Please HELP! Can you guys show your steps too? Thankss :}

OpenStudy (anonymous):

Create a proportion :)

OpenStudy (anonymous):

Don't chu gotta use : I= p*r*t ?

OpenStudy (anonymous):

umm...

OpenStudy (anonymous):

Lol..sorry l learned this but i forgot...sorry :(

OpenStudy (anonymous):

Aww :[ Okay thanks anyway :\

OpenStudy (anonymous):

No problem sista :) if you need help with anything else other then commision i will be here to help :)

OpenStudy (anonymous):

Use interest=P*R*t I=650*(2/100)*5=65 Net amount=P+I=650+65=715